State and prove Pascal’s identity using the formula for "{n \choose k}"


1
Expert's answer
2021-11-08T15:36:40-0500

Pascal's identity states that "{n \choose k} = {n-1 \choose k-1}+{n-1 \choose k}", where "{n \choose k}={\frac{n!} {(n-k)!*k!}}"

"{n-1 \choose k-1}+{n-1 \choose k}= {\frac{(n-1)!} {(n-1-k+1)!*(k-1)!}}+{\frac{(n-1)!} {(n-1-k)!*k!}}={\frac{(n-1)!} {(n-k)!*(k-1)!}}+{\frac{(n-1)!} {(n-k-1)!*k!}}={\frac{k*(n-1)!+(n-k)*(n-1)!} {(n-k)!*k!}}={\frac{k*(n-1)!+n*(n-1)!-k*(n-1)!} {(n-k)!*k!}}={\frac{n!} {(n-k)!*k!}}"

The statement has been proven